Let f(x)=x−8x+8
f
(
x
)
=
x

8
x
+
8
. Find the open intervals on which f
f
is increasing (decreasing). Then determine the x
x
-coordinates of all relative maxima (minima).
1
Expert's answer
2020-04-30T10:23:22-0400

For finding the intervals where function is increasing or decreasing

we first differentiate the function and we get f'(x),


and for increasing interval we solve the inequality f'(x)>0,

and for decreasing interval we solve the inequality f'(x)<0


and for finding the maxima or minima we solve for f'(x)=0 and analyze further.

We have "f(x)= \n\\frac{\nx\u22128}{x+8}=1-\\frac{16}{x+8}\n\u200b" ,

differentiating we get

"f'(x)=\frac{16}{(x+8)^2}"

so we can see that "f'(x)" is a squared quantity which will be positive over all the interval, hence "f(x)" is an increasing function

and "f'(x)" has no roots so the maximum and minimum will be positive and negative infinity.

I have tried to show this function by a graph attached below.